40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Mar 5 11:33:10 2014, Koji, Summary, LSC, 2 arm ALS->LSC transition - road map 
    Reply  Wed Mar 5 19:15:39 2014, Jenne, Summary, LSC, ALS offset moving script modified 
    Reply  Wed Mar 5 22:32:21 2014, manasa, Update, LSC, Stuck at step 2 
       Reply  Thu Mar 6 11:15:32 2014, Koji, Summary, LSC, Stuck at step 2 
    Reply  Fri Mar 7 00:43:34 2014, manasa, Update, LSC, ALS C&D locked (on MC2 and ETMs) and IR resonance obtained ALS_MC2CARM.pdf
Message ID: 9696     Entry time: Wed Mar 5 22:32:21 2014     In reply to: 9691     Reply to this: 9698
Author: manasa 
Type: Update 
Category: LSC 
Subject: Stuck at step 2 

Quote:

Step by step description of transition from 2arm ALS to Common/Differential LSC for FPMI

- Step 0: Place the frequencies of the arm green beams at the opposite side of the carrier green.

- Step 1: Activate stablization loops for ALSX and ALSY simultaneously.
  (Use LSC filter modules for the control. This still requires correct handling of the servo and filter module triggers)

- Step 2: Activate stablization loops for ALS Common and Differential by actuating ETMX and ETMY

I locked the arms using ALS error signals and the LSC filter modules. But when I try to acquire CARM and DARM using ALS, the arms lose lock when the matrix elements ALSX to Yarm and ALSY to X arm reach -/+0.9

What I did:

1. ALS locking of arms
(i) Found arm beat notes
(ii) Input matrix POX and POY elements set to '0'
(iii) Aux matrix elements ALSX to Xarm and ALSY to Y arm set to '1'
(iv) Power normalization matrix elements for TRX and TRY set to '0'
(v) Triggers for arm lock over ridden and the FM triggers were set to 'manual'
(vi) Arm servo gains set to '0'
(vii) All but FM5 were disabled
(viii) Phase tracker history reset and servo actuation set to ETMs
(ix) Servo gain increased in steps (+/-10 for the arms)
(x) FM1, FM6, FM7 enabled (see note 1 below)
(xi) FM9 enabled

Arms were locked with ~2000Hz rms

2. CARM and DARM locking
(i) Scanned the arms for IR resonance
(ii) Moved off-resonance (Stepped arm servo offsets by 30 counts)
(iiI) Stepped matrix elements ALSY to X arm and ALSX to Y arm ezcastep C1:LSC-PD_DOF_MTRX_6_29 +-0.1 C1:LSC-PD_DOF_MTRX_7_28 +0.1

Whenever the matrix elements reached -/+0.9, the arms were kicked out of lock. I don't see anything obvious as to why this is happening even after nearly 10+times of redoing.

Notes:
1. I found the filters for the arm servos different for X and Y. FM1 and FM8 were missing in one of the filter modules. Jenne remembered Rana modifying and removing the unnecessary filters in one arm. We put back FM1 (low pass filter) which might not be necessary for PDH lock but is necessary for ALS. FM8 is now added to FM7.
2. To self : Check ALS Y arm power outlets (60Hz frequency comb seen in the error signal)

ELOG V3.1.3-